OpenStudy (anonymous):

What must be added to 3x -7 to make x2 + 4x -1?

OpenStudy (anonymous):

x x^2 and 6

OpenStudy (anonymous):

3x+x =4x x^2=x^2 -7+6=-1

OpenStudy (anonymous):

x^2 + 4x + -1

OpenStudy (anonymous):

or -1

OpenStudy (anonymous):

thank you ^.^